OpenStudy (anonymous):

An experiment consists of rolling two fair dice and adding the dots on the two sides facing up. What is the probability that the sum of the dots is 5?

OpenStudy (anonymous):

5=2+3 5=3+2 5=1+4 5=4+1 There are 4 possibilities out of 36 4/36=1/9

OpenStudy (anonymous):

thanks, that's what i got also.

OpenStudy (anonymous):

Great

OpenStudy (anonymous):

what about this one?A fair cube with faces numbered from 1 through 6 is tossed a number of times. If the probability of the cube landing on the number 2 every time is 1/216, what is the best answer for how many times the cube is rolled?